OpenStudy (anonymous):

You run a canoe-rental business on a small river in Ohio. You currently charge $12 per canoe and average 36 rentals a day. An industry journal says that, for every fifty-cent increase in rental price, the average business can expect to lose two rentals a day. Use this information to attempt to maximize your income. What should you charge?

OpenStudy (anonymous):

If the price per rental = 12 + x(0.50) then number of rentals = 36 -(x*2) total income = (12+.5x) (36-2x) the equation will be -x^2 - 6x +432 the max income will occur at -b/2a of a quadratic equation = -3 and its value will be -D/4a = 441

OpenStudy (anonymous):

thanks...!! :D